OpenStudy (anonymous):

Solve for x: (x + 4) / 2 = 7. x = 3 x = 5 x = 6 x = 10

OpenStudy (anonymous):

x=10

OpenStudy (texaschic101):

(x + 4) / 2 = 7 --- multiply both sides by 2 x + 4 = 14 -- subtract 4 from both sides x = 14 - 4 x = 10

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Solve for x: 1/3 (2x - 8) = 4. (1 point) 2 6 10 10/3

OpenStudy (anonymous):

@texaschic101